OpenStudy (anonymous):

find f(10) when f(x)= -2x+9

OpenStudy (anonymous):

-11 right?

jimthompson5910 (jim_thompson5910):

f(x) = -2x + 9 f(10) = -2(10) + 9 ... replace every x with 10 f(10) = -20 + 9 f(10) = -11 So you are 100% correct

OpenStudy (anonymous):

ok, how about this one...

OpenStudy (anonymous):

find f(-x) when f(x) =4x+3

OpenStudy (anonymous):

-4x-3?

OpenStudy (anonymous):

it has to be written into an equation

jimthompson5910 (jim_thompson5910):

no, f(10) means you replace every x with 10 in f(x) f(-x) means you replace every 'x' with '-x'

OpenStudy (anonymous):

but how do i write it into an equation

jimthompson5910 (jim_thompson5910):

f(x) =4x+3 f(-x) =4(-x)+3 ... replace every x with -x and simplify

OpenStudy (anonymous):

so would it be -4x+3?

OpenStudy (anonymous):

i understand the replacing.. but the equation, i JUST replace and write it as -4x+3?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

f(-x) = -4x + 3

OpenStudy (anonymous):

ok,,, thank u! :)

jimthompson5910 (jim_thompson5910):

np